OpenStudy (anonymous):

please integrate the e^(2x) cos3x dx

OpenStudy (tommo_lcfc):

This uses the product rule, if you know that.

OpenStudy (tommo_lcfc):

I'm just doing the question.

OpenStudy (tommo_lcfc):

sorry, it says integrate, hang on

OpenStudy (tommo_lcfc):

Integrating by parts

OpenStudy (dumbcow):

you will have to do integration by parts twice

OpenStudy (tommo_lcfc):

Or use the shortcut method.

OpenStudy (anonymous):

please can you show me how to do it?

OpenStudy (anonymous):

@tommo_lcfc , is that the tabular method you're referring to?

OpenStudy (tommo_lcfc):

Tell me what it is and I'll tell you if I'm referring to it or not. I was taught this method at university.

OpenStudy (anonymous):

@tommo_lcfc..what kind of method is that?

OpenStudy (tommo_lcfc):

Basically, if f(x) is a polynomial, we just differentiate this until it becomes constant and integrate the other term.

OpenStudy (tommo_lcfc):

I understand that the question does not refer to a polynomial but the trig differentiation is repeated. You always get either sin, cos, -sin or -cos (unless you are differentiating something else).

OpenStudy (tommo_lcfc):

Let's see what you all get.

OpenStudy (anonymous):

Let u = e^2x and dv = cos(3x)dx. Then, du = 2e^2x dx and v = sin(3x)/3 The integral may be rewritten as follows ∫u dv = uv - ∫v du ∫ e^(2x) cos(3x) dx = e^(2x) sin(3x)/3 - (2/3)∫e^(2x) sin(3x) dx T his requires us to do integration by parts again p = e^(2x), dp = 2e^(2x) dx, dq = sin(3x) dx, q = -cos(3x)/3 Now, ∫p dq = pq - ∫q dp ∫ e^(2x) cos(3x) dx = e^(2x) sin(3x)/3 - (2/3)[-e^(2x) cos(3x)/3 + (2/3)∫e^(2x) cos(3x) dx] = e^(2x) sin(3x)/3 + (2/9)e^(2x) cos(3x) - (4/9)∫e^(2x) cos(3x) dx You have 9/9 of the original integral on the LHS and -4/9 of it on the RHS. So add 4/9 to both sides (13/9)∫ e^(2x) cos(3x) dx = e^(2x) sin(3x)/3 + (2/9)e^(2x) cos(3x) Finally, multiply both sides by 9/13 and add the constant of integration to get the final answer ∫ e^(2x) cos(3x) dx = (3/13) e^(2x) sin(3x) + (2/13) e^(2x) cos(3x) + C this is their answer but i was confused in this equation: ∫ e^(2x) cos(3x) dx = e^(2x) sin(3x)/3 - (2/3)[-e^(2x) cos(3x)/3 + (2/3)∫e^(2x) cos(3x) dx] where did he get the sin(3x)/3 please can you explain it to me?

OpenStudy (tommo_lcfc):

sin(3x)/3 is what you get when you integrate sin(3x) as you divide the component inside the x term, in this case 3.

OpenStudy (tommo_lcfc):

Also, I didn't get that answer. I got 3e^(2x) [5/4 sin(3x) - 15/8 cos(3x)] + c using my method.

OpenStudy (lalaly):

\[\huge{\int\limits{e^{2x}\cos(3x)dx}}\] let u=cos(3x) dv=e^2xdx du=-3sin(3x)dx v=e^2x/2 \[\int\limits{e^{2x}\cos(3x)}=\cos(3x) \frac{e^{2x}}{2}-\int\limits{\frac{e^{2x}}{2}\times (-3\sin(3x))dx}\]\[\int\limits{e^{2x}\cos(3x)dx}=\frac{1}{2}\cos(3x)e^{2x}+\frac{3}{2}\int\limits{e^{2x}\sin(3x)dx}\] now u do the second integral by parts \[\int\limits{e^{2x}\sin(3x)dx}\] let u=sin(3x) dv=e^2xdx du=3cos(3x)dx v=e^2x/2 so\[\int\limits{e^{2x}\sin(3x)dx}=\sin(3x)\frac{e^{2x}}{2}-\frac{3}{2} \int\limits{e^{2x}\cos(3x)dx}\] now substitute this in the second equation i wrote\[\int\limits{e^{2x}\cos(3x)dx}=\frac{1}{2}\cos(3x)e^{2x}+\frac{3}{2}[\frac{1}{2}\sin(3x)e^{2x}-\frac{3}{2}\int\limits{e^{2x}\cos(3x)dx)}\]now u have the same integral on both sides .. move one on the other side and solve for the integral

OpenStudy (experimentx):

i usually take it as formula http://answers.yahoo.com/question/index?qid=20080723161756AA9xWm6

OpenStudy (lalaly):

thats a good idea :D but also everyone needs to know how to do it in case they forgot the formula :D:D

OpenStudy (experimentx):

yeah ... it's the general rule :D same goes to sin bx :D